OpenStudy (anonymous):

The perimeter of a triangle is 70cm. The longest side is 4cm less than the sum of the other two sides. Twice the shortest side is 9cm less than the longest side. Find the length of each side of the triangle.

OpenStudy (shamil98):

Let's get some variables in here and make an equation.. You have a given perimeter = 70. So P = 70. Longest side is L = ( x + y) - 4 where x and y are the other sides. 2x = L - 9

OpenStudy (shamil98):

2x+9 = L we need to find y then substitute L = x + y - 4 -y = - L + x -4 y = L - x + 4

OpenStudy (shamil98):

L + x + y = 70

OpenStudy (shamil98):

2x + 9 + x + 2x + 9 - x + 4 = 70

OpenStudy (shamil98):

4x + 22 = 70 4x = 48 x = 12

OpenStudy (shamil98):

Okay, so now you have the value of x figured out, do you know how to get the value of L and y?

OpenStudy (shamil98):

L + 12 + y = 70 L + y = 58 Remember 2x + 9 = L so solve for L now.

OpenStudy (shamil98):

if x = 12, and 2x + 9 = L L = ?

OpenStudy (anonymous):

I'm still trying to process how you found x. It's so different compared to how I thought it was suppose to be worked out.

OpenStudy (e.mccormick):

There are different ways... all depends on how you want to do it. L = S + M - 4 L + S + M = 70 2S = L-9 L - S - M = - 4 L + S + M = 70 -L + 2S+0M =-9 Can be soved as a matrix or three equation system, but what shamil98 is showing works great too.

OpenStudy (shamil98):

Alright. The perimeter of a triangle is 70cm. The longest side is 4cm less than the sum of the other two sides. Twice the shortest side is 9cm less than the longest side P = 70 Longest side is 4 less than the sum of the other sides is Longest side = (x+y) - 4 or L = (x+y) - 4 twice the shortest side is 9 less than the longest side so 2x = L - 9 2x + 9 = L you have two equations now. you want to find out what y is equal to and then substitute y and L and solve for x. which is what I did. L = x + y - 4 -L - y -L - y -y = -L + x - 4 divide by -1 y = L - x + 4 so y = L - x + 4 L = 2x + 9 and L + x +y = 70 let's plug in those values into L + x + y = 70 2x + 9 + x + (L - x + 4) = 70 3x + 9 + 2x + 9 - x + 4 = 70 combine the like terms 4x + 22 = 70 4x = 48 x = 12 Follow so far?..

OpenStudy (anonymous):

I somewhat see how you did that. Are the two equations: L=(x+y)-4 and 2x+9=L?

OpenStudy (anonymous):

I think I understand how to get y now: substitute the 12 in for x in the equation: 2x=(x+y-4)-9, which gives me 25.

OpenStudy (shamil98):

Yep, so know you want to get the Longest side.

OpenStudy (shamil98):

L = 2x + 9

OpenStudy (anonymous):

How do I get the longest side?

OpenStudy (shamil98):

x = 12, L = 2x + 9

OpenStudy (anonymous):

Ok, I see! Thank you so much!

OpenStudy (e.mccormick):

So, you got them all now?

OpenStudy (e.mccormick):

I mentioned the matrix method of solving this. In case you wondered what I meant: L = S + M - 4 L + S + M = 70 2S = L-9 L - S - M = - 4 L + S + M = 70 -L + 2S+0M =-9 So I have three rows, R1, R2, and R3. If I use elementary row operations to get them to pivots, a value of 1 on the left of the | in each row and column, it goes like this: \(\left[\begin{array}{ccc|c} 1 & -1 & -1 & -4\\ 1 & 1 & 1 & 70\\ -1 & 2 & 0 & -9 \end{array}\right]\) Add R1 and R2 new R1 \(\left[\begin{array}{ccc|c} 2 & 0 & 0 & 66\\ 1 & 1 & 1 & 70\\ -1 & 2 & 0 & -9 \end{array}\right]\) R1 \(\times \dfrac{1}{2}\) \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 1 & 1 & 1 & 70\\ -1 & 2 & 0 & -9 \end{array}\right]\) R2 + (-1)R1 new R2 R3 + R1 new R1 \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 0 & 1 & 1 & 37\\ 0 & 2 & 0 & 24 \end{array}\right]\) R3 \(\times \dfrac{1}{2}\) \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 0 & 1 & 1 & 37\\ 0 & 1 & 0 & 12 \end{array}\right]\) R2 + (-1)R3 \(\left[\begin{array}{ccc|c} 1 & 0 & 0 & 33\\ 0 & 0 & 1 & 25\\ 0 & 1 & 0 & 12 \end{array}\right]\) Longest = 33cm Medium = 25cm Shortest = 12cm And a reference on this: http://www.epcc.edu/tutorialservices/valleverde/Documents/Gauss-Jordan_Method.pdf

OpenStudy (anonymous):

Yes, I have all of the values now. Thank you, I appreciate you showing me the alternate way on how to do it.

OpenStudy (e.mccormick):

np. have fun!